# OPERA: A Reinforcement Learning–Enhanced Orchestrated Planner-Executor Architecture for Reasoning-Oriented Multi-Hop Retrieval

### v1

@@ Abstract
Recent advances in large language models (LLMs) and dense retrievers have driven significant progress in retrieval-augmented generation (RAG). However, existing approaches face significant challenges in complex reasoning-oriented multi-hop retrieval tasks: 1) Ineffective reasoning-oriented planning: Prior methods struggle to generate robust multi-step plans for complex queries, as rule-based decomposers perform poorly on out-of-template questions. 2) Suboptimal reasoning-driven retrieval: Related methods employ limited query reformulation, leading to iterative retrieval loops that often fail to locate golden documents. 3) Insufficient reasoning-guided filtering: Prevailing methods lack the fine-grained reasoning to effectively filter salient information from noisy results, hindering utilization of retrieved knowledge. Fundamentally, these limitations all stem from the weak coupling between retrieval and reasoning in current RAG architectures. We introduce the Orchestrated Planner-Executor Reasoning Architecture (OPERA), a novel reasoning-driven retrieval framework. OPERA’s Goal Planning Module (GPM) decomposes questions into sub-goals, which are executed by a Reason-Execute Module (REM) with specialized components for precise reasoning and effective retrieval. To train OPERA, we propose Multi-Agents Progressive Group Relative Policy Optimization (MAPGRPO), a novel variant of GRPO. Experiments on complex multi-hop benchmarks show OPERA’s superior performance, validating both the MAPGRPO method and OPERA’s design. Code is available at https://github.com/Ameame1/OPERA.
{strip}[Uncaptioned image]
Figure 1: Overview of OPERA’s MAPGRPO training framework and performance comparison with traditional RAG.
@@ 1 Introduction
The ability to solve complex problems is a core aspect of intelligence, and within Retrieval-Augmented Generation (RAG), reasoning-centric retrieval provides an effective means to address these tasks in the post-training era. The concurrent improvement of Large Language Models (LLMs) (Brown et al. 2020; Devlin et al. 2019) and dense retrieval systems (Karpukhin et al. 2020; Khattab and Zaharia 2020) has propelled the evolution of RAG (Lewis et al. 2020). The traditional RAG follows a retrieve-then-reason paradigm (Lee, Chang, and Toutanova 2019; Lewis et al. 2020), which has now been widely optimized into a multi-stage pipeline including query rewriting, document retrieval, document filtering, and answer generation (Izacard and Grave 2021). However, despite significant progress, effectively orchestrating these capabilities remains challenging, as existing approaches struggle with the demands of multi-hop reasoning. Consider a query such as, “What was the previous occupation of the person who succeeded the founder of the company that acquired WhatsApp?” Such questions demand not merely retrieving documents, but orchestrating a precise sequence of retrieval and reasoning steps where each operation depends critically on its predecessors.
Current approaches face several key challenges. First, existing solutions have limited reasoning-oriented planning. While planner-first models like PlanRAG (Lee, An, and Kim 2024) and REAPER (Joshi et al. 2024) introduce upfront planning mechanisms, their static plans cannot dynamically adapt to unforeseen challenges during retrieval. Second, most methods lack effective reasoning-aware retrieval. Even adaptive methods like Adaptive-RAG (Jeong et al. 2024) and AT-RAG (Rezaei et al. 2024), which adjust retrieval strategy based on query complexity, lack the fine-grained, reasoning-driven query reformulation needed for complex multi-hop scenarios. Third, current systems provide inadequate reasoning-guided filtering. Even when retrieval fetches golden documents, they are often buried within noisy top-K results. While iterative approaches like ReAct (Yao et al. 2023b) attempt to address this through reasoning-action loops, and recent methods such as BGE (Ke et al. 2024) show promise in bridging retriever-LLM preferences, their effectiveness remains limited by indirect reward signals that fail to capture the nuanced reasoning required for effective filtering. These limitations persist because advanced enhancement strategies—spanning SFT, preference optimization, and RL—are insufficient or misaligned, often leading to goal misalignment between modules (Qi et al. 2024; Wang et al. 2023). These issues stem from fundamental weakness in the coupling between retrieval and reasoning, preventing full utilization of modern LLMs and dense retrievers’ capabilities.
@@
  • Strong empirical results on multi-hop benchmarks. Extensive experiments validate OPERA’s reasoning-centric architecture and the effectiveness of our training approach.
@@ 2 Related Work
RAG. To mitigate hallucination, Retrieval-Augmented Generation (RAG) was introduced to ground outputs in external knowledge (Lewis et al. 2020). The initial “retrieve-then-read” paradigm (Lee, Chang, and Toutanova 2019; Lewis et al. 2020), however, proved insufficient for multi-hop reasoning due to its static pipeline (Trivedi et al. 2022; Arabzadeh, Yan, and Clarke 2021; Luan et al. 2021). The field has since evolved toward more dynamic retrieval, for instance by routing queries based on complexity (Jeong et al. 2024) or introducing explicit planning (Lee, An, and Kim 2024). While these methods improve upon static RAG, they primarily optimize the retrieval act itself, rather than the overarching reasoning strategy. OPERA is distinct in its hierarchical approach: a specialized Goal Planning Module (GPM) governs high-level strategy, while an agentic Reason-Execute Module (REM) handles tactical execution, including fine-grained analysis and adaptive query reformulation.
@@
Chain-of-Thought. Chain-of-Thought (CoT) methods (Wei et al. 2022; Yao et al. 2023a) and agentic frameworks like ReAct (Yao et al. 2023b) and IRCoT (Trivedi et al. 2023) established the importance of decomposing problems and interleaving reasoning with actions. However, these frameworks rely on a single, general LLM for high-level planning and low-level execution, which can compromise reliability. While multi-agent systems like MetaGPT (Hong et al. 2024) assign distinct roles, they typically use generalist models. OPERA advances by employing an asymmetric architecture, pairing the strategic Goal Planning Module (GPM) with the agentic Reason-Execute Module (REM) to separate strategic and tactical concerns.
@@
Reinforcement Learning. A key aspect of our work is the training of our specialized planner and executor. Many frameworks utilize on-policy algorithms like Proximal Policy Optimization (PPO); however, PPO often struggles with the large action spaces and sparse rewards common in RAG (Stiennon et al. 2020; Uc-Cetina et al. 2022; Ramamurthy et al. 2023), making training unstable and sample-inefficient. To address this, preference-based optimization has gained traction. While Direct Preference Optimization (DPO) (Rafailov et al. 2023) has become a standard for learning from binary preferences (chosen >> rejected), it is ill-suited for more nuanced reward signals (Ivison et al. 2024). Our training methodology generates fine-grained scalar scores reflecting the quality of a plan or an execution step. Using DPO would discard this rich information by compressing it into a binary signal. To fully leverage these scalar rewards within our multi-agent setting, we build upon Group Relative Policy Optimization (GRPO) (Shao et al. 2024) and introduce our own variant: Multi-Agents Progressive Group Relative Policy Optimization (MAPGRPO). Unlike standard GRPO, MAPGRPO is specifically designed for our staged training protocol, enabling fine-grained credit assignment (Papangelis et al. 2019) and ensuring coordinated optimization across the distinct roles of the GPM and REM agents.
@@ 3 Method
We introduce OPERA (Orchestrated Planner-Executor Reasoning Architecture), a framework that systematically decouples strategic planning from tactical execution. As illustrated in Figures 1 and 2, OPERA operates through two core modules: the Goal Planning Module (GPM) containing a Plan Agent for strategic decomposition, and the Reason-Execute Module (REM) containing Analysis-Answer and Rewrite Agents for conditional execution and adaptive retrieval. The Plan Agent decomposes complex questions into sub-goals $\mathcal{P}$ (plan consisting of sub-goals) with placeholder dependencies. The Analysis-Answer Agent performs information sufficiency assessment $\phi$ (information sufficiency indicator) and answer extraction from retrieved documents $\mathcal{D}_{i}$ (documents retrieved for sub-goal $i$). The Rewrite Agent reformulates queries when information is insufficient. To optimize this multi-agent system, we introduce MAPGRPO for sequential training with role-specific rewards.
@@
We propose MAPGRPO, a novel variant of Group Relative Policy Optimization (GRPO) (Shao et al. 2024).
@@ 4 Experimental
Datasets. We evaluate OPERA on three multi-hop reasoning benchmarks: HotpotQA (Yang et al. 2018) (90K questions), 2WikiMultiHopQA (Ho et al. 2020) (150K questions), and Musique (Trivedi et al. 2022) (25K questions). For out-of-domain evaluation, we use NQ (Kwiatkowski et al. 2019) and MultiHopRAG (Tang and Yang 2024).
Implementation Details. We use Qwen2.5-7B-Instruct***https://huggingface.co/Qwen/Qwen2.5-7B-Instruct for Plan and Analysis-Answer agents and other baseline’s backbone, Qwen2.5-3B-Instructhttps://huggingface.co/Qwen/Qwen2.5-3B-Instruct for the Rewrite agent, and BGE-M3 (Chen et al. 2024) as our dense retriever with top-5 document retrieval. For pre-scored dataset construction, we utilize the DeepSeek R1 APIDeepSeek R1 is accessed via the official API service provided by DeepSeek (https://platform.deepseek.com). as the data generation model. Training details are in Appendix A.2.
Baselines. We compare OPERA against methods in four categories: Naive: (1) Qwen2.5-7B (No Retrieval); (2) Single-Step RAG. CoT: (3) IRCoT (Trivedi et al. 2023); (4) OPERA (CoT Only). SFT: (5) Adaptive-RAG (Jeong et al. 2024). RL: (6) BGM (Ke et al. 2024); (7) OPERA (MAPGRPO). All SFT and RL methods are trained on mixed datasets, with Plan Agent utilizing pre-scored dataset $\mathcal{D}_{\text{scored}}$ for high-score sample selection (Appendix A.6). Baselines use defaults with Faiss (Douze et al. 2024).
@@
Adaptive-RAG (Jeong et al. 2024) 45.7 56.9 30.1 39.3 24.3 35.7 SFT
BGM (Ke et al. 2024) 41.5 53.8 44.3 55.8 19.6 26.8 RL
@@
These results indicate that OPERA’s performance gains stem from the synergistic combination of architectural design and coordinated training: while specialized agents with clearly defined responsibilities provide the foundational framework, the coordinated optimization through MAPGRPO ensures effective collaboration across the planning, retrieval, and reasoning workflow.
@@
Runtime Dynamics and Attention Flow. Figure 3 illustrates OPERA’s decision-making and execution patterns. The left panel shows agent activation over time: the Plan Agent (blue) initiates question cycles, the Analysis-Answer Agent (green) performs reasoning, and the Rewrite Agent (orange) activates upon retrieval failures. The black trajectory tracks the cumulative question completion rate, demonstrating performance across three questions of varying complexity. The right panel illustrates attention evolution across processing stages—from entity-focused planning to relation-aware analysis and context-integrated refinement.
Stable Convergence and Reward Evolution. MAPGRPO enables stable and efficient training with consistent performance improvements. The training dynamics are shown in Figures 4 and 5. As shown in Figure 4, training exhibits typical RL characteristics: initial instability (epochs 1-5), followed by progressive improvement with occasional performance dips, particularly visible in the Rewrite Agent due to its conditional activation. Crucially, our expert demonstration strategy (Expert Injection Impact) proves highly effective at reducing policy gradient variance, a key factor in training stability. The shaded region in the bottom-right panel highlights the Expert Injection Impact, showing a significant variance reduction for the 7B model with expert injection compared to its counterpart without, confirming our theoretical analysis.
@@
Latency Analysis Across Questions. Due to latency variations across questions of different complexity, we analyze 100 real-world multi-hop test questions to examine agent call patterns and individual component contributions. Figure 6 shows that the Plan Agent maintains relatively consistent latency, while the Analysis-Answer Agent shows higher variance depending on reasoning complexity. The Rewrite Agent activates sporadically when retrieval failures occur, confirming OPERA’s adaptive behavior and practical efficiency for real-world deployment.
Out-of-Domain Evaluation. Table 3 shows how training methods affect generalization. On NQ—single-hop QA over full Wikipedia pages where planning is unnecessary—MAPGRPO achieves 36.6% EM while SFT drops to 19.5% (from 23.1% CoT baseline). MAPGRPO preserves OPERA’s architectural flexibility, allowing it to bypass planning for single-hop queries and use the Analysis-Answer Agent’s training on (sub-question, document, answer) tuples to handle long documents. SFT appears to overfit to multi-hop patterns, attempting decomposition even when not needed. On MHRAG, which has multi-hop structure similar to our training data, all methods improve, with MAPGRPO reaching 55.7% EM. RL methods perform well on both single and multi-hop tasks while SFT performs worse on single-hop tasks. This suggests that trajectory-based optimization with diverse rewards helps OPERA adapt its reasoning strategy to different tasks, while supervised learning tends toward consistent multi-hop behavior.
@@ 5 Conclusion
We propose OPERA, a multi-agent framework that addresses limitations in RAG systems through specialized planning and execution roles. OPERA shows improvements—reaching 39.7% EM on Musique (63.4% relative improvement) and exceeding 60% EM on 2WikiMultiHopQA—by combining architectural design and MAPGRPO training, where specialized agents provide separation of concerns while role-specific rewards enable coordination. Ablation studies show that architectural components have impact on performance than training methods, with Plan Agent removal causing drops below untrained baselines. This suggests that reasoning-driven retrieval benefits from architectural changes beyond optimization. OPERA has limitations with ambiguous decomposition paths and 4+ hop questions, but generalization to out-of-domain tasks indicates the approach works across reasoning patterns. Future work could explore dynamic agent selection, incorporate commonsense knowledge, and extend OPERA’s principles to other reasoning tasks.

### v2

@@ Abstract
Recent advances in large language models (LLMs) and dense retrievers have driven significant progress in retrieval-augmented generation (RAG). However, existing approaches face significant challenges in complex reasoning-oriented multi-hop retrieval tasks: 1) Ineffective reasoning-oriented planning: Prior methods struggle to generate robust multi-step plans for complex queries, as rule-based decomposers perform poorly on out-of-template questions. 2) Suboptimal reasoning-driven retrieval: Related methods employ limited query reformulation, leading to iterative retrieval loops that often fail to locate golden documents. 3) Insufficient reasoning-guided filtering: Prevailing methods lack the fine-grained reasoning to effectively filter salient information from noisy results, hindering utilization of retrieved knowledge. Fundamentally, these limitations all stem from the weak coupling between retrieval and reasoning in current RAG architectures. We introduce the Orchestrated Planner-Executor Reasoning Architecture (OPERA), a novel reasoning-driven retrieval framework. OPERA’s Goal Planning Module (GPM) decomposes questions into sub-goals, which are executed by a Reason-Execute Module (REM) with specialized components for precise reasoning and effective retrieval. To train OPERA, we propose Multi-Agents Progressive Group Relative Policy Optimization (MAPGRPO), a novel variant of GRPO. Experiments on complex multi-hop benchmarks show OPERA’s superior performance, validating both the MAPGRPO method and OPERA’s design.
Codehttps://github.com/Ameame1/OPERA
Extended versionhttps://arxiv.org/abs/2508.16438
@@ 1 Introduction
The ability to solve complex problems is a core aspect of intelligence, and within Retrieval-Augmented Generation (RAG), reasoning-centric retrieval provides an effective means to address these tasks in the post-training era. The concurrent improvement of Large Language Models (LLMs) (brown2020language; devlin2019bert) and dense retrieval systems (karpukhin2020dense; khattab2020colbert) has propelled the evolution of RAG. The traditional RAG follows a retrieve-then-reason paradigm (lee2019latent; lewis2020retrieval), which has now been widely optimized into a multi-stage pipeline including query rewriting, document retrieval, document filtering, and answer generation (izacard2020leveraging). However, despite significant progress, effectively orchestrating these capabilities remains challenging, as existing approaches struggle with the demands of multi-hop reasoning. Consider a query such as, “What was the previous occupation of the person who succeeded the founder of the company that acquired WhatsApp?” Such questions demand not merely retrieving documents, but orchestrating a precise sequence of retrieval and reasoning steps where each operation depends critically on its predecessors.
Current approaches face several key challenges. First, existing solutions have limited reasoning-oriented planning. While planner-first models like PlanRAG (lee2024planrag) and REAPER (reaper2024) introduce upfront planning mechanisms, their static plans cannot dynamically adapt to unforeseen challenges during retrieval. Second, most methods lack effective reasoning-aware retrieval. Even adaptive methods like Adaptive-RAG (jeong2024adaptive) and AT-RAG (zhang2024atrag), which adjust retrieval strategy based on query complexity, lack the fine-grained, reasoning-driven query reformulation needed for complex multi-hop scenarios. Third, current systems provide inadequate reasoning-guided filtering. Even when retrieval fetches golden documents, they are often buried within noisy top-K results. While iterative approaches like ReAct (yao2022react) attempt to address this through reasoning-action loops, and recent methods such as BGE (ke2024bridging) show promise in bridging retriever-LLM preferences, their effectiveness remains limited by indirect reward signals that fail to capture the nuanced reasoning required for effective filtering. These limitations persist because advanced enhancement strategies—spanning SFT, preference optimization, and RL—are insufficient or misaligned, often leading to goal misalignment between modules (qi2023fine; wang2023instruction). These issues stem from fundamental weakness in the coupling between retrieval and reasoning, preventing full utilization of modern LLMs and dense retrievers’ capabilities.
@@
  • Strong empirical results on multi-hop benchmarks. Extensive experiments validate OPERA’s reasoning-centric architecture and training approach are effective.
@@ 2 Related Work
RAG. To mitigate hallucination, Retrieval-Augmented Generation (RAG) was introduced to ground outputs in external knowledge (lewis2020retrieval). The initial “retrieve-then-read” paradigm (lee2019latent), however, proved insufficient for multi-hop reasoning due to its static pipeline (trivedi2022musique; arabzadeh2021predicting; luan2021sparse). The field has since evolved toward more dynamic retrieval, for instance by routing queries based on complexity (jeong2024adaptive) or introducing explicit planning (lee2024planrag). While these methods improve upon static RAG, they primarily optimize the retrieval act itself, rather than the overarching reasoning strategy. OPERA is distinct in its hierarchical approach: a specialized Goal Planning Module (GPM) governs high-level strategy, while an agentic Reason-Execute Module (REM) handles tactical execution, including fine-grained analysis and adaptive query reformulation.
@@
Chain-of-Thought. Chain-of-Thought (CoT) methods (wei2022chain; yao2023tree) and agentic frameworks like ReAct (yao2022react) and IRCoT (trivedi2023interleaving) established the importance of decomposing problems and interleaving reasoning with actions. However, these frameworks rely on a single, general LLM for high-level planning and low-level execution, which can compromise reliability. While multi-agent systems like MetaGPT (hong2024metagpt) assign distinct roles, they typically use generalist models. OPERA advances by employing an asymmetric architecture, pairing the strategic Goal Planning Module (GPM) with the agentic Reason-Execute Module (REM) to separate strategic and tactical concerns.
@@
Reinforcement Learning. A key aspect of our work is the training of our specialized planner and executor. Many frameworks utilize on-policy algorithms like Proximal Policy Optimization (PPO); however, PPO often struggles with the large action spaces and sparse rewards common in RAG (stiennon2020learning; fernandezoLooria2022survey; ramamurthy2023reinforcement), making training unstable and sample-inefficient. To address this, preference-based optimization has gained traction. While Direct Preference Optimization (DPO) (rafailov2023direct) has become a standard for learning from binary preferences (chosen >> rejected), it is ill-suited for more nuanced reward signals (ivison2024unpacking). Our training methodology generates fine-grained scalar scores reflecting the quality of a plan or an execution step. Using DPO would discard this rich information by compressing it into a binary signal. To fully leverage these scalar rewards within our multi-agent setting, we build upon Group Relative Policy Optimization (GRPO) (shao2024grpo) and introduce our own variant: Multi-Agents Progressive Group Relative Policy Optimization (MAPGRPO). Unlike standard GRPO, MAPGRPO is specifically designed for our staged training protocol, enabling fine-grained credit assignment (papangelis2019collaborative) and ensuring coordinated optimization across the distinct roles of the GPM and REM agents.
@@ 3 Method
We introduce OPERA (Orchestrated Planner-Executor Reasoning Architecture), a framework that systematically decouples strategic planning from tactical execution. As illustrated in Figures LABEL:fig:motivation and 2, OPERA operates through two core modules: the Goal Planning Module (GPM) containing a Plan Agent for strategic decomposition, and the Reason-Execute Module (REM) containing Analysis-Answer and Rewrite Agents for conditional execution and adaptive retrieval. The Plan Agent decomposes complex questions into sub-goals $\mathcal{P}$ (plan consisting of sub-goals) with placeholder dependencies. The Analysis-Answer Agent performs information sufficiency assessment $\phi$ (information sufficiency indicator) and answer extraction from retrieved documents $\mathcal{D}_{i}$ (documents retrieved for sub-goal $i$). The Rewrite Agent reformulates queries when information is insufficient. To optimize this multi-agent system, we introduce MAPGRPO for sequential training with role-specific rewards.
@@
We propose MAPGRPO, a novel variant of Group Relative Policy Optimization (GRPO) (shao2024grpo).
@@ 4 Experimental
Datasets. We evaluate OPERA on three multi-hop reasoning benchmarks: HotpotQA (yang2018hotpotqa) (90K questions), 2WikiMultiHopQA (ho2020constructing) (150K questions), and Musique (trivedi2022musique) (25K questions). For out-of-domain evaluation, we use NQ (kwiatkowski2019natural) and MultiHopRAG (tang2024multihop).
Implementation Details. We use Qwen2.5-7B-Instruct***https://huggingface.co/Qwen/Qwen2.5-7B-Instruct for Plan and Analysis-Answer agents and other baseline’s backbone, Qwen2.5-3B-Instructhttps://huggingface.co/Qwen/Qwen2.5-3B-Instruct for the Rewrite agent, and BGE-M3 (chen2024bge) as our dense retriever with top-5 document retrieval. For pre-scored dataset construction, we utilize the DeepSeek R1 APIDeepSeek R1 is accessed via the official API service provided by DeepSeek (https://platform.deepseek.com). as the data generation model. Training details are in Appendix A.2.
Baselines. We compare OPERA against methods in four categories: Naive: (1) Qwen2.5-7B (No Retrieval); (2) Single-Step RAG. CoT: (3) IRCoT (trivedi2023interleaving); (4) OPERA (CoT Only). SFT: (5) Adaptive-RAG (jeong2024adaptive). RL: (6) BGM (ke2024bridging); (7) OPERA (MAPGRPO). All SFT and RL methods are trained on mixed datasets, with Plan Agent utilizing pre-scored dataset $\mathcal{D}_{\text{scored}}$ for high-score sample selection (Appendix A.6). Baselines use defaults with Faiss (douze2024faiss).
@@
Adaptive-RAG (jeong2024adaptive) 45.7 56.9 30.1 39.3 24.3 35.7 SFT
BGM (ke2024bridging) 41.5 53.8 44.3 55.8 19.6 26.8 RL
@@
These results indicate that OPERA’s performance gains stem from the synergistic blend of architectural design and training: while specialized agents with defined responsibilities provide the foundational framework, coordinated optimization through MAPGRPO ensures effective collaboration in planning, retrieval, and reasoning workflow.
@@
Runtime Dynamics and Attention Flow. Figure 3 shows OPERA’s decision-making and execution patterns. The left panel shows agent activation over time: the Plan Agent (blue) initiates question cycles, the Analysis-Answer Agent (green) performs reasoning, and the Rewrite Agent (orange) activates upon retrieval failures. The black trajectory tracks the cumulative question completion rate, demonstrating performance across three questions of varying complexity. The right panel illustrates attention evolution across processing stages—from entity-focused planning to relation-aware analysis and context-integrated refinement.
Stable Convergence and Reward Evolution. MAPGRPO enables stable and efficient training with consistent performance improvements. Training dynamics are shown in Figures 4 and 5. Figure 4 shows typical RL dynamics: initial instability (epochs 1–5) followed by gradual improvement with occasional dips, most evident in the conditionally activated Rewrite Agent. The shaded region in the bottom-right panel highlights the Expert Injection Impact, showing significant variance reduction for the 7B model with expert injection compared to its counterpart without, effectively improving training stability and confirming our theoretical analysis.
@@
Latency Analysis Across Questions. To analyze latency variations across questions of varying complexity, we evaluate 100 multi-hop test questions, focusing on agent call patterns and component contributions. Figure 6 shows that the Plan Agent maintains relatively consistent latency, while the Analysis-Answer Agent shows higher variance depending on reasoning complexity. The Rewrite Agent activates only when retrieval failures occur, confirming OPERA’s adaptive behavior and efficiency for deployment.
Out-of-Domain Evaluation. Table 3 shows how training methods affect generalization. On NQ—single-hop QA over full Wikipedia pages where planning is unnecessary—MAPGRPO achieves 36.6% EM while SFT drops to 19.5% (from 23.1% CoT baseline). MAPGRPO preserves OPERA’s flexibility, allowing it to bypass planning for single-hop queries and use the Analysis-Answer Agent’s training on (sub-question, document, answer) tuples to handle long documents. SFT appears to overfit to multi-hop patterns, attempting decomposition even when not needed. On MHRAG, which has multi-hop structure similar to our training data, all methods improve, with MAPGRPO reaching 55.7% EM. RL methods perform well on both single and multi-hop tasks while SFT performs worse on single-hop tasks. This suggests that trajectory-based optimization with diverse rewards helps OPERA adapt its reasoning strategy to different tasks, while supervised learning tends toward consistent multi-hop behavior.
@@ 5 Conclusion
We propose OPERA, a multi-agent framework that addresses limitations in RAG systems through specialized planning and execution roles. OPERA shows improvements—reaching 39.7% EM on Musique (63.4% relative improvement) and exceeding 60% EM on 2WikiMultiHopQA—by combining architectural design and MAPGRPO training, where specialized agents provide separation of concerns while role-specific rewards enable coordination. Ablation studies show that architectural design impacts performance more than training, with Plan Agent removal dropping below untrained baselines. This suggests that reasoning-driven retrieval benefits from architectural advances beyond optimization. While OPERA still struggles with ambiguous decomposition and long reasoning chains, its generalization to out-of-domain tasks demonstrates robustness across reasoning patterns.